# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 3 Existence of $(r,\delta)$-surgical solutions
From now on we fix $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7, and the constant $\varepsilon_{4}$ is as in Proposition 2.5. Choose $0<\varepsilon<\frac{1}{4}\min\{\frac{1}{10^{3}n},\tilde{\varepsilon}_{1}(\varepsilon_{0})\}$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6, and $\theta>0$. Let $\beta=\beta(\varepsilon)$ be the constant given in Lemma 5.8, and choose
@@
(cf. [6])    Let $n\geq 5$, $0<\varepsilon^{\prime}<\varepsilon_{1}$, where $\varepsilon_{1}$ is from Proposition 5.4, $f$ and $\theta$ in the pinching assumption be fixed, and $A$ and $C^{\prime}$ be two positive numbers. Then there exist $Q>0$ and $\Lambda>0$ with the following property. Assume that $(\mathcal{O}(t),g(t))$, $t\in[a,b]$, is a complete surgical solution (with at most isolated orbifold singularities) to the $n$-dimensional Ricci flow with positive isotropic curvature and $(f,\theta)$-pinched, bounded curvature. Let $(x_{0},t_{0})$ be a space-time point (in $\mathcal{O}(t_{0})$) which is not in a component with strictly PIC2, such that:
@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Then $0<\rho\leq A$. It is easy to see that there exists, up to an extraction, $\bar{y}_{k}\in B(\bar{x}_{k},0,\rho)$ such that
@@
We pull back the rescaled solutions $(U_{k},\bar{g}_{k}(\cdot))$ to $\widetilde{U}_{k}$ via the covering maps $\pi_{k}:\widetilde{U}_{k}\rightarrow U_{k}$. Choose a point $\tilde{z}_{k}\in\pi_{k}^{-1}(\bar{z}_{k})$, and let $\tilde{\gamma}_{k}$ be the component of $\pi_{k}^{-1}(\bar{\gamma}_{k})$ which contains the point $\tilde{z}_{k}$. By using a local compactness theorem for the Ricci flow (see Theorem C.3.3 in [6], cf. also Theorem 4.1.5 in [27], Theorem 16.1 in [47], and Appendix E in [55]) we see that up to an extraction, the sequence of pointed solutions $(B(\tilde{z}_{k},0,\frac{\rho-a}{2}),\pi_{k}^{*}\bar{g}_{k}(\cdot),(\tilde{z}_{k},0))$ (where $B(\tilde{z}_{k},0,\frac{\rho-a}{2})=B_{\pi_{k}^{*}\bar{g}_{k}(0)}(\tilde{z}_{k},\frac{\rho-a}{2}))$ converges smoothly to a partial Ricci flow $(B(z_{\infty},0,\frac{\rho-a}{2}),g_{\infty}(\cdot),(z_{\infty},0))$, and the sequence of geodesics $\tilde{\gamma}_{k}$ converges to a (minimal) geodesic $\gamma_{\infty}$ (w.r.t. $g_{\infty}(0)$) starting from $z_{\infty}$. The geodesic $\gamma_{\infty}$ is covered by strong $2\varepsilon^{\prime}$-necks, and the scalar curvature goes to infinity as a point approaches the open end of the image of $\gamma_{\infty}$. By Lemma 5.5, $B(z_{\infty},0,\frac{\rho-a}{2})$ contains a $2\varepsilon^{\prime}$-horn, say $H$, diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, which contains the open end of the image of $\gamma_{\infty}$. For any $z\in B(z_{\infty},0,\frac{\rho-a}{2})$, the evolving metric $g_{\infty}(z,\cdot)$ exists at least on the time interval $[-\frac{R(z,0)^{-1}}{2},0]$. In addition, by using the pinching assumption we see that $(B(z_{\infty},0,\frac{\rho-a}{2}),g_{\infty}(\cdot))$ is weakly PIC2 and strictly PIC.
Let $(\overline{B}_{\infty},\hat{d})$ be the completion of the metric space $(B(z_{\infty},0,\frac{\rho-a}{2}),d_{\infty})$, where the metric $d_{\infty}$ is induced by the Riemannian metric $g_{\infty}(0)$. Let $y_{\infty}$ be the limit point of $\gamma_{\infty}$ which lies in $\overline{B}_{\infty}\setminus B(z_{\infty},0,\frac{\rho-a}{2})$, and $\widehat{U}_{\infty}=H\cup\{y_{\infty}\}$. The pointed metric space $(\widehat{U}_{\infty},\hat{d},y_{\infty})$ is a locally complete Alexandrov space with nonnegative curvature, and there exists a $n$-dimensional tangent cone of $(\widehat{U}_{\infty},\hat{d})$ at $y_{\infty}$ with small aperture, denoted by $C_{\infty}$, which is a nonnegatively curved metric cone. Denote the vertex of the cone $C_{\infty}$ by 0. Then a sequence of suitable rescalings of certain pieces of $(H,g_{\infty}(\cdot))$ converges smoothly to a Ricci flow of weakly PIC2 and strict PIC whose final time slice lies in $C_{\infty}\setminus\{0\}$. (For more details, see Step 3 in the proof of Theorem 6.1.1 in [6] and Step 2 in the proof of Theorem 4.1 in [31]; compare also [20], [56] and [71].) This contradicts Proposition 6.5 in [20]. $\Box$
@@
(cf. [7])    Let $n\geq 5$, $f$ and $\theta$ in the pinching assumption be fixed. Let $\varepsilon>0$ and $C>0$ be chosen as above. Given $r,\delta>0$, there exist $h\in(0,\delta r)$ and $D>10$, such that if $(\mathcal{O}(\cdot),g(\cdot))$ is a complete surgical solution to the Ricci flow in dimension $n$ with positive isotropic curvature and with bounded curvature, defined on a time interval $[0,b]$ and satisfying the $(f,\theta)$-pinching assumption and the canonical neighborhood assumption $(CN)_{r}$ with $(4\varepsilon,4C)$-control, then the following holds:
@@
Proof.  We essentially follow the proof of Theorem 6.2.1 in [6], Theorem 6.5 in [7], and Proposition 11.1 in [20] with some necessary modifications. (Compare the proof of Propositions 4.2 and 4.4 in [29], and Proposition 4.2 in [51].) We argue by contradiction. Otherwise, there exist $r,\delta>0$, sequences $h_{k}\rightarrow 0$, $D_{k}\rightarrow+\infty$, a sequence of complete surgical solutions $(\mathcal{O}_{k}(\cdot),g_{k}(\cdot))$ in dimension $n$ with positive isotropic curvature and bounded curvature and satisfying the $(f,\theta)$-pinching assumption and $(CN)_{r}$ with $(4\varepsilon,4C)$-control, and sequences $t_{k}>0$, $x_{k},y_{k},z_{k}\in\mathcal{O}_{k}(t_{k})$ with $R(x_{k},t_{k})\leq 2r^{-2}$, $R(y_{k},t_{k})=h_{k}^{-2}$ and $R(z_{k},t_{k})\geq D_{k}h_{k}^{-2}$, and finally a sequence of curves $\gamma_{k}$ in $\mathcal{O}_{k}(t_{k})$ connecting $x_{k}$ to $z_{k}$ and containing $y_{k}$, whose points of scalar curvature in $[8Cr^{-2},(4C)^{-1}D_{k}h_{k}^{-2}]$ are centers of $4\varepsilon$-necks, but $(y_{k},t_{k})$ is not a center of a strong $\delta$-neck.
@@
Now we describe Hamilton’s metric surgery on a $\delta$-neck $N$ in a Riemannian orbifold $(\mathcal{O},g)$ with a diffeomorphism $\psi:\mathbb{S}^{n-1}/\Gamma\times(-\frac{1}{\delta},\frac{1}{\delta})\rightarrow N$ which is a $\delta$-homothety with scaling factor $Q$. Roughly speaking, the surgery is removing the open subset $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},(10\delta)^{-1}))$ from $\mathcal{O}$ and gluing in a Riemannian orbifold (called a surgery cap) diffeomorphic to $D^{n}//\Gamma$ (which has an isolated singular point when $\Gamma$ is nontrivial) along each boundary component thus produced. More precisely, the gluing at the boundary component $\psi(\mathbb{S}^{n-1}/\Gamma\times\{-(10\delta)^{-1}\})$ of $\mathcal{O}\setminus\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},(10\delta)^{-1}))$ is as follows; the gluing at $\psi(\mathbb{S}^{n-1}/\Gamma\times\{(10\delta)^{-1}\})$ is similar. As in the appendix and the proof of Proposition 2.5, we choose a smooth function $\varphi:(-(10\delta)^{-1},B-(10\delta)^{-1}]\rightarrow\mathbb{R}$, where $B>\frac{1}{10}$, (in particular, $\varphi(z)=ce^{-1/(z+(10\delta)^{-1})}$ when $z\in(-(10\delta)^{-1},\frac{1}{10}-(10\delta)^{-1}]$, where $0<c<\frac{1}{\sqrt{(n-1)(n-2)}}$,) and a smooth cutoff function $\chi:(\frac{1}{20}-(10\delta)^{-1},B-(10\delta)^{-1}]\rightarrow[0,1]$ with $\chi(z)=1$ for $\frac{1}{20}-(10\delta)^{-1}<z\leq\frac{1}{18}-(10\delta)^{-1}$ and $\chi(z)=0$ for $\frac{1}{12}-(10\delta)^{-1}\leq z\leq B-(10\delta)^{-1}$, and deform the metric $g$ on $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},B-(10\delta)^{-1}])$ to a new metric $g^{\prime}$, with
$g^{\prime}(x)=\begin{cases}g(x),&{z\in(-\frac{1}{\delta},-\frac{1}{10\delta}];}\\ e^{-2\varphi(z)}g(x),&{z\in(-\frac{1}{10\delta},\frac{1}{20}-\frac{1}{10\delta}];}\\ e^{-2\varphi(z)}(\chi(z)g(x)+(1-\chi(z))Q^{-1}(\psi_{*}\bar{g}_{\text{cyl}})(x)),&{z\in(\frac{1}{20}-\frac{1}{10\delta},B-\frac{1}{10\delta}],}\end{cases}$
@@
where for $x\in\psi(\mathbb{S}^{n-1}/\Gamma\times(-\frac{1}{\delta},B-\frac{1}{10\delta}])$ we denote $z(x)=p_{2}\circ\psi^{-1}(x)$ by $z$, here $p_{2}:\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},\delta^{-1})\rightarrow(-\delta^{-1},\delta^{-1})$ is the projection onto the second factor, and $\bar{g}_{\text{cyl}}$ is the standard metric on the (quotient) cylinder $\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},\delta^{-1})$ with scalar curvature 1. (We require that $\frac{1}{\delta}\gg B$.) In particular, the function $\varphi$ is chosen such that the metric $g^{\prime}$ restricted to the portion $\psi(\mathbb{S}^{n-1}/\Gamma\times[B^{\prime}-(10\delta)^{-1},B-(10\delta)^{-1}])$ has constant positive sectional curvature, here $\frac{1}{10}<B^{\prime}<B$ as in the appendix. Then we smoothly attach a quotient of a metric $n$-ball of constant positive sectional curvature by $\Gamma$ to $\mathcal{O}\setminus\psi(\mathbb{S}^{n-1}/\Gamma\times(B-(10\delta)^{-1},(10\delta)^{-1}-B))$ (with the new metric) along the boundary component $\psi(\mathbb{S}^{n-1}/\Gamma\times\{B-(10\delta)^{-1}\})$, as we do in the construction of the orbifold standard solution in the appendix. Note that there is a positive number $\delta^{\prime}=\delta^{\prime}(\delta)$ with $\lim_{\delta\rightarrow 0}\delta^{\prime}=0$ such that the union of $\psi(\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},-(10\delta)^{-1}])$ and the surgery cap with the metric just described is $\delta^{\prime}$-homothetic to the union of $\mathbb{S}^{n-1}/\Gamma\times(-\frac{9}{10\delta},0]$ and the surgery cap with the metric $\hat{g}_{\Gamma}(0)$ defined in the appendix. (Here, $\delta^{\prime}$ also depends on the choice of the functions $\varphi$ and $\chi$ above.)
@@
Definition (compare [7], [51]).   Given an interval $I\subset[0,+\infty)$, fix surgery parameters $r$, $\delta>0$ and let $h$, $D$, $\Theta=2Dh^{-2}$ be the associated cutoff parameters. Let $(\mathcal{O}(t),g(t))$ ($t\in I$) be an evolving compact Riemannian orbifold with at most isolated singularities. Let $t_{0}\in I$ and $(\mathcal{O}_{+},g_{+})$ be a (possibly empty) Riemmanian $n$-orbifold. We say that $(\mathcal{O}_{+},g_{+})$ is obtained from $(\mathcal{O}(\cdot),g(\cdot))$ by $(r,\delta)$-surgery at time $t_{0}$ if
i. $R_{\text{max}}(g(t_{0}))=\Theta$, and there is a finite collection of pairwise disjoint strong $\delta$-necks in $\mathcal{O}(t_{0})$, centered at some points with scalar curvature equal to $h^{-2}$ at time $t_{0}$, such that $\mathcal{O}_{+}$ is obtained from $\mathcal{O}(t_{0})$ by doing Hamilton’s surgery along these necks as described above, and removing each of the following components:
@@
Let $n\geq 12$, the family $\mathcal{F}_{t}$, $t\in[0,\hat{T}]$, be as above, and $f$ and $\theta$ be as in (3.1). Let $\varepsilon$ and $C$ be chosen as before. Let $\delta$ be sufficiently small such that (an orbifold version of) Proposition 8.2 in [20] holds. Let $(\mathcal{O}(t),g(t))$, $t\in[0,b]$, be a compact $(r,\delta)$-surgical solution to the Ricci flow in dimension $n$ (with $(f,\theta)$-pinched curvature), where $b<\hat{T}$. Suppose that the curvature tensor at time $b$ lies in $\mathcal{F}_{b}$, and $R_{max}(b)=\Theta$, which is defined above. Then there exists a Riemannian orbifold $(\mathcal{O}_{+},g_{+})$ which is obtained from $(\mathcal{O}(\cdot),g(\cdot))$ by $(r,\delta)$-surgery at time $b$, such that the curvature tensor of $g_{+}$ still lies in the set $\mathcal{F}_{b}$, and $R_{min}(g_{+})\geq R_{min}(g(b))$.

### v22

@@ 3 Existence of $(r,\delta)$-surgical solutions
From now on we fix $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition A.7, and the constant $\varepsilon_{4}$ is as in Proposition 2.5. Choose $0<\varepsilon<\frac{1}{4}\min\{\frac{1}{10^{4}n},\tilde{\varepsilon}_{1}(\varepsilon_{0})\}$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma A.6, and $\theta>0$. Let $\beta=\beta(\varepsilon)$ be the constant given in Lemma A.8, and choose
@@
(cf. [6])    Let $n\geq 5$, $0<\varepsilon^{\prime}<\varepsilon_{1}$, where $\varepsilon_{1}$ is from Proposition A.4, $f$ and $\theta$ in the pinching assumption be fixed, and $A$ and $C^{\prime}$ be two positive numbers. Then there exist numbers $Q>0$ and $\Lambda>0$ with the following property. Assume that $(\mathcal{O}(t),g(t))$, $t\in[a,b]$, is a complete surgical solution (with at most isolated orbifold singularities) to the $n$-dimensional Ricci flow with positive isotropic curvature and $(f,\theta)$-pinched, bounded curvature. Let $(x_{0},t_{0})$ be a space-time point (in $\mathcal{O}(t_{0})$) which is not in a compact component with strictly PIC2, such that:
@@ 1. $R(x_{k},t_{k})\geq Q_{k}$;
Then $0<\rho\leq A$. It is easy to see that there exists, up to an extraction, a sequence of points $\bar{y}_{k}\in B(\bar{x}_{k},0,\rho)$ such that
@@
We pull back the rescaled solutions $(U_{k},\bar{g}_{k}(\cdot))$ to $\widetilde{U}_{k}$ via the covering maps $\pi_{k}:\widetilde{U}_{k}\rightarrow U_{k}$. Choose a point $\tilde{z}_{k}\in\pi_{k}^{-1}(\bar{z}_{k})$, and let $\tilde{\gamma}_{k}$ be the component of $\pi_{k}^{-1}(\bar{\gamma}_{k})$ which contains the point $\tilde{z}_{k}$. By using a local compactness theorem for the Ricci flow (see Theorem C.3.3 in [6], cf. also Theorem 4.1.5 in [29], Theorem 16.1 in [50], and Appendix E in [58]) we see that up to an extraction, the sequence of pointed solutions $(B(\tilde{z}_{k},0,\frac{\rho-a}{2}),\pi_{k}^{*}\bar{g}_{k}(\cdot),(\tilde{z}_{k},0))$ (where $B(\tilde{z}_{k},0,\frac{\rho-a}{2})=B_{\pi_{k}^{*}\bar{g}_{k}(0)}(\tilde{z}_{k},\frac{\rho-a}{2}))$ converges smoothly to a partial Ricci flow $(B(z_{\infty},0,\frac{\rho-a}{2}),g_{\infty}(\cdot),(z_{\infty},0))$, and the sequence of geodesics $\tilde{\gamma}_{k}$ converges to a (minimal) geodesic $\gamma_{\infty}$ (w.r.t. $g_{\infty}(0)$) starting from $z_{\infty}$. The geodesic $\gamma_{\infty}$ is covered by strong $2\varepsilon^{\prime}$-necks, and the scalar curvature goes to infinity as a point approaches the open end of the image of $\gamma_{\infty}$. By Lemma A.5, $B(z_{\infty},0,\frac{\rho-a}{2})$ contains a $2\varepsilon^{\prime}$-horn, say $H$, diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, which contains the open end of the image of $\gamma_{\infty}$. For any $z\in B(z_{\infty},0,\frac{\rho-a}{2})$, the evolving metric $g_{\infty}(z,\cdot)$ exists at least on the time interval $[-\frac{R(z,0)^{-1}}{2},0]$. In addition, by using the pinching assumption we see that $(B(z_{\infty},0,\frac{\rho-a}{2}),g_{\infty}(\cdot))$ is weakly PIC2 and strictly PIC.
Let $(\overline{B}_{\infty},\hat{d})$ be the completion of the metric space $(B(z_{\infty},0,\frac{\rho-a}{2}),d_{\infty})$, where the metric $d_{\infty}$ is induced by the Riemannian metric $g_{\infty}(0)$. Let $y_{\infty}$ be the limit point of $\gamma_{\infty}$ which lies in $\overline{B}_{\infty}\setminus B(z_{\infty},0,\frac{\rho-a}{2})$, and $\widehat{U}_{\infty}=H\cup\{y_{\infty}\}$. The pointed metric space $(\widehat{U}_{\infty},\hat{d},y_{\infty})$ is a locally complete Alexandrov space with nonnegative curvature, and there exists an $n$-dimensional tangent cone of $(\widehat{U}_{\infty},\hat{d})$ at $y_{\infty}$ with small aperture, denoted by $C_{\infty}$, which is a nonnegatively curved metric cone. Denote the vertex of the cone $C_{\infty}$ by 0. Then a sequence of suitable rescalings of certain pieces of $(H,g_{\infty}(\cdot))$ converges smoothly to a Ricci flow of weakly PIC2 and strict PIC whose final time slice lies in $C_{\infty}\setminus\{0\}$. (For more details, see Step 3 in the proof of Theorem 6.1.1 in [6] and Step 2 in the proof of Theorem 4.1 in [33]; compare also [21], [59] and [75].) This contradicts Proposition 6.5 in [21]. $\Box$
@@
(cf. [7])    Let $n\geq 5$, $f$ and $\theta$ in the pinching assumption be fixed. Let $\varepsilon>0$ and $C>0$ be chosen as above. Given $r,\delta>0$, there exist numbers $h\in(0,\delta r)$ and $D>10$, such that if $(\mathcal{O}(\cdot),g(\cdot))$ is a complete surgical solution to the Ricci flow in dimension $n$ with positive isotropic curvature and with bounded curvature, defined on a time interval $[0,b]$ and satisfying the $(f,\theta)$-pinching assumption and the canonical neighborhood assumption $(CN)_{r}$ with $(4\varepsilon,4C)$-control, then the following holds:
@@
Proof.  We essentially follow the proof of Theorem 6.2.1 in [6], Theorem 6.5 in [7], and Proposition 11.1 in [21] with some necessary modifications. (Compare the proof of Propositions 4.2 and 4.4 in [31], and Proposition 4.2 in [54].) We argue by contradiction. Otherwise, there exist numbers $r,\delta>0$, sequences $h_{k}\rightarrow 0$, $D_{k}\rightarrow+\infty$, a sequence of complete surgical solutions $(\mathcal{O}_{k}(\cdot),g_{k}(\cdot))$ in dimension $n$ with positive isotropic curvature and bounded curvature and satisfying the $(f,\theta)$-pinching assumption and $(CN)_{r}$ with $(4\varepsilon,4C)$-control, and sequences $t_{k}>0$, $x_{k},y_{k},z_{k}\in\mathcal{O}_{k}(t_{k})$ with $R(x_{k},t_{k})\leq 2r^{-2}$, $R(y_{k},t_{k})=h_{k}^{-2}$ and $R(z_{k},t_{k})\geq D_{k}h_{k}^{-2}$, and finally a sequence of curves $\gamma_{k}$ in $\mathcal{O}_{k}(t_{k})$ connecting $x_{k}$ to $z_{k}$ and containing $y_{k}$, whose points of scalar curvature in $[8Cr^{-2},(4C)^{-1}D_{k}h_{k}^{-2}]$ are centers of $4\varepsilon$-necks, but $(y_{k},t_{k})$ is not a center of a strong $\delta$-neck.
@@
Now we describe Hamilton’s metric surgery on a $\delta$-neck $N$ (with $\delta$ small) in a Riemannian orbifold $(\mathcal{O},g)$ with a diffeomorphism $\psi:\mathbb{S}^{n-1}/\Gamma\times(-\frac{1}{\delta},\frac{1}{\delta})\rightarrow N$ which is a $\delta$-homothety with scaling factor $Q$. Roughly speaking, the surgery is removing the open subset $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},(10\delta)^{-1}))$ from $\mathcal{O}$ and gluing in a Riemannian orbifold (called a surgery cap) diffeomorphic to $D^{n}//\Gamma$ (which has an isolated singular point when $\Gamma$ is nontrivial) along each boundary component thus produced. More precisely, the gluing at the boundary component $\psi(\mathbb{S}^{n-1}/\Gamma\times\{-(10\delta)^{-1}\})$ of $\mathcal{O}\setminus\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},(10\delta)^{-1}))$ is as follows; the gluing at $\psi(\mathbb{S}^{n-1}/\Gamma\times\{(10\delta)^{-1}\})$ is similar. As in the appendix and the proof of Proposition 2.5, we choose a smooth function $\varphi:(-(10\delta)^{-1},b_{0}-(10\delta)^{-1}]\rightarrow\mathbb{R}$, where $b_{0}>\frac{1}{10}$, (in particular, $\varphi(z)=ce^{-1/(z+(10\delta)^{-1})}$ when $z\in(-(10\delta)^{-1},\frac{1}{10}-(10\delta)^{-1}]$, where $0<c<\frac{1}{\sqrt{(n-1)(n-2)}}$ is a small positive constant,) and a smooth cutoff function $\chi:(\frac{1}{20}-(10\delta)^{-1},b_{0}-(10\delta)^{-1}]\rightarrow[0,1]$ with $\chi(z)=1$ for $\frac{1}{20}-(10\delta)^{-1}<z\leq\frac{1}{18}-(10\delta)^{-1}$ and $\chi(z)=0$ for $\frac{1}{12}-(10\delta)^{-1}\leq z\leq b_{0}-(10\delta)^{-1}$, and deform the metric $g$ on $\psi(\mathbb{S}^{n-1}/\Gamma\times(-(10\delta)^{-1},b_{0}-(10\delta)^{-1}])$ to a new metric $g^{\prime}$, with
$g^{\prime}(x)=\begin{cases}g(x),&{z\in(-\frac{1}{\delta},-\frac{1}{10\delta}];}\\ e^{-2\varphi(z)}g(x),&{z\in(-\frac{1}{10\delta},\frac{1}{20}-\frac{1}{10\delta}];}\\ e^{-2\varphi(z)}(\chi(z)g(x)+(1-\chi(z))Q^{-1}(\psi_{*}\bar{g}_{\text{cyl}})(x)),&{z\in(\frac{1}{20}-\frac{1}{10\delta},b_{0}-\frac{1}{10\delta}],}\end{cases}$
@@
where for $x\in\psi(\mathbb{S}^{n-1}/\Gamma\times(-\frac{1}{\delta},b_{0}-\frac{1}{10\delta}])$ we denote $z(x)=p_{2}\circ\psi^{-1}(x)$ by $z$, here $p_{2}:\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},\delta^{-1})\rightarrow(-\delta^{-1},\delta^{-1})$ is the projection onto the second factor, and $\bar{g}_{\text{cyl}}$ is the standard metric on the (quotient) cylinder $\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},\delta^{-1})$ with scalar curvature 1. (We require that $\frac{1}{\delta}\gg b_{0}$.) In particular, the function $\varphi$ is chosen such that the metric $g^{\prime}$ restricted to the portion $\psi(\mathbb{S}^{n-1}/\Gamma\times[b^{\prime}-(10\delta)^{-1},b_{0}-(10\delta)^{-1}])$ has constant positive sectional curvature, here $\frac{1}{10}<b^{\prime}<b_{0}$ as in the appendix. Then we smoothly attach a quotient of a closed metric $n$-ball of constant positive sectional curvature by $\Gamma$ to $\mathcal{O}\setminus\psi(\mathbb{S}^{n-1}/\Gamma\times(b_{0}-(10\delta)^{-1},(10\delta)^{-1}-b_{0}))$ (with the new metric) along the boundary component $\psi(\mathbb{S}^{n-1}/\Gamma\times\{b_{0}-(10\delta)^{-1}\})$, as we do in the construction of the orbifold standard solution in the appendix. Note that there is a positive number $\delta^{\prime}=\delta^{\prime}(\delta)$ with $\lim_{\delta\rightarrow 0}\delta^{\prime}=0$ such that the union of $\psi(\mathbb{S}^{n-1}/\Gamma\times(-\delta^{-1},-(10\delta)^{-1}])$ and the surgery cap with the metric just described is $\delta^{\prime}$-homothetic to the union of $\mathbb{S}^{n-1}/\Gamma\times(-\frac{9}{10\delta},0]$ and the surgery cap with the metric $\hat{g}_{\Gamma}(0)$ defined in the appendix. (Here, $\delta^{\prime}$ also depends on the choice of the functions $\varphi$ and $\chi$ above.)
@@
Definition (compare [7], [54]).   Given an interval $I\subset[0,+\infty)$, fix surgery parameters $r$, $\delta>0$ and let $h$, $D$, $\Theta=2Dh^{-2}$ be the associated cutoff parameters. Let $(\mathcal{O}(t),g(t))$ ($t\in I$) be an evolving complete Riemannian orbifold with at most isolated singularities. Let $t_{0}\in I$ and $(\mathcal{O}_{+},g_{+})$ be a (possibly empty) Riemmanian $n$-orbifold. We say that $(\mathcal{O}_{+},g_{+})$ is obtained from $(\mathcal{O}(\cdot),g(\cdot))$ by $(r,\delta)$-surgery at time $t_{0}$ if
i. $R_{\text{max}}(g(t_{0}))=\Theta$, and there is a collection of pairwise disjoint strong $\delta$-necks in $\mathcal{O}(t_{0})$, centered at some points with scalar curvature equal to $h^{-2}$ at time $t_{0}$, such that $\mathcal{O}_{+}$ is obtained from $\mathcal{O}(t_{0})$ by doing Hamilton’s surgery along these necks as described above, and removing each of the following components:
@@
Let $n\geq 12$, the family $\mathcal{F}_{t}$, $t\in[0,\hat{T}]$, be as above, and $f$ and $\theta$ be as in (3.1). Let $\varepsilon$ and $C$ be chosen as before. Let $\delta$ be sufficiently small such that $\delta\ll\varepsilon$ and (an orbifold version of) Proposition 8.2 in [21] holds. Let $(\mathcal{O}(t),g(t))$, $t\in[0,b]$, be a compact $(r,\delta)$-surgical solution to the Ricci flow of dimension $n$ (with $(f,\theta)$-pinched curvature), where $b<\hat{T}$. Suppose that the curvature tensor at time $b$ lies in $\mathcal{F}_{b}$, and $R_{max}(b)=\Theta$, which is defined above. Then there exists a Riemannian orbifold $(\mathcal{O}_{+},g_{+})$ which is obtained from $(\mathcal{O}(\cdot),g(\cdot))$ by $(r,\delta)$-surgery at time $b$, such that the curvature tensor of $g_{+}$ still lies in the set $\mathcal{F}_{b}$, and $R_{min}(g_{+})\geq R_{min}(g(b))$.